40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 4 11:10:28 2010, nancy, Update, IOO, Mode Cleaner WFS 
    Reply  Wed Aug 4 12:49:42 2010, nancy, Update, IOO, Mode Cleaner WFS 6x
       Reply  Thu Aug 5 01:29:39 2010, nancy, Update, IOO, Mode Cleaner WFS 
          Reply  Thu Aug 5 11:48:52 2010, nancy, Update, IOO, Mode Cleaner WFS 
             Reply  Thu Aug 5 13:05:53 2010, Koji, Update, IOO, Mode Cleaner WFS 
                Reply  Fri Aug 6 10:44:55 2010, nancy, koji, Update, IOO, Mode Cleaner WFS 
                   Reply  Fri Aug 6 12:33:10 2010, Koji, Update, IOO, Mode Cleaner WFS 
                      Reply  Mon Aug 9 12:46:24 2010, Nancy, Update, IOO, Mode Cleaner WFS 
Message ID: 3372     Entry time: Fri Aug 6 10:44:55 2010     In reply to: 3367     Reply to this: 3374
Author: nancy, koji 
Type: Update 
Category: IOO 
Subject: Mode Cleaner WFS 

Quote:

Upon Nancy's request, I checked the status of the suspensions.

I found that the power strip of the 1Y4 rack was turned off.
Since it has a over current breaker, I don't know whether it happened by someone or over current.

Anyway, I restarted the sus computers, and now the suspensions are damping as usual.
The MC has been aligned, the auto locker is also working.


Incidentally, I found that the WFS servos are not working. Actually since the last night
It repeated losing lock and unlock.

Probably some values of the matrix or the gain is wrong.
I left the WFS as it is because Nancy will put new values this afternoon.
I will ask her to confirm that the old values work at the end of her work.

 Yesterday , I put in the Output Matrix, and changed the gain sliders for the 4 WFS loops.

It worked and was keeping the lock for the MC.

I then tested whether the MC1 and 3 were following any change in MC2 alignment. It was indeed workinng,

Next we stepped to putting in the gains for the MC2 oplev servo.

the signs are decided on the basis of convergence, and the magnitude is kept very low, to have a very slow control for MC2.

This complete 6 * 6 model does work, and was able to keep the transmission held.

I also tried poking each mirror in pitcg and yaw, and the cavity comes back to high resonance after some time.

This time is indeed large if the poking is made for MC2, and the transmission comes back to normal after big oscillations.

I tried to measure the Open loop TFs for all these loops yesterday, but somehow could not find a correct excitation.

I will do it today.

Plan ahead :

1.  Center the spot on MC2 and the QPD

2. Optimize the gains by looking at response to noise.

3. Measure Power Spectrum Density of each error signal.

 

ELOG V3.1.3-